# Online POMDP Planning with Anytime Deterministic Guarantees

### v1

@@ 3 Mathematical Analysis
where the simplified expectation operator, $\bar{\mathbb{E}}[\cdot]$, is taken with respect to the unnormalized likelihood ¯$\bar{\mathbb{P}}(z_{t+1}\mid H^{-}_{t+1})$.
3.1 Simplified observation space
We first analyze the performance guarantees of a simplified observation space, while assuming a complete belief update at each belief state, i.e., $\bar{\mathcal{X}}\equiv\mathcal{X}$. The following theorem describes the guarantees of the observation-simplified value function with respect to its theoretical value,
Theorem 1
@@ 3.1 Simplified observation space
$\left|V^{\pi}(b_{t})\!-\!\bar{V}^{\pi}(b_{t})\right|\leq\!\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\!\sum_{z_{t+1:{\tau}}}\sum_{x_{t:{\tau}}}b(x_{t})\!\!\prod_{k=t+1}^{\tau}\overline{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]\triangleq\epsilon_{z}^{\pi}(b_{t}),$ (9)
where we use a subscript of $(z)$ in $\epsilon_{z}^{\pi}(b_{t})$ to denote observation-only simplification. Importantly, the bound only contains terms which depend on observations that are within the simplified space, $z\in\bar{\mathcal{Z}}$. This is an essential property of the bound, as it is a value that can easily be calculated during the planning process and provides a certification of the policy quality at any given node along the tree. Furthermore, it is apparent from (9) that as the number of observations included in the simplified set, $\bar{\mathcal{Z}}$, increases, the value of $\epsilon_{z}^{\pi}(b_{t})$ consequently diminishes,
leading to a convergence towards the theoretical value function, i.e. $\epsilon_{z}^{\pi}(b_{t})\rightarrow 0$.
Theorem 1 provides both lower and upper bounds for the theoretical value function, assuming a fixed policy. Using this theorem, we can derive upper and lower bounds for any policy, including the optimal one. This is achieved by applying the Bellman optimality operator to the upper bound in a repeated manner, instead of the estimated value function; In the context of tree search algorithms, our algorithm explores only a subset of the decision tree due to pruned observations. However, at every belief node encountered during this exploration, all potential actions are expanded. The action-value of these expanded actions is bounded using the Upper Deterministic Bound, which we now define as
$\textsc{UDB}^{\pi}(b_{t},a_{t})\triangleq\bar{Q}^{\pi}(b_{t},a_{t})+\epsilon_{z}^{\pi}(b_{t},a_{t})=r(b_{t},a_{t})+\bar{\mathbb{E}}_{z_{t+1}}[\bar{V}^{\pi}(b_{t+1})]+\epsilon_{z}^{\pi}(b_{t},a_{t}),$ (10)
where the action-dependent bound on the value difference, $\epsilon_{z}^{\pi}(b_{t},a_{t})$, is the bound of taking action $a_{t}$ in belief $b_{t}$ and following policy $\pi$ thereafter,
$\epsilon_{z}^{\pi}(b_{t},a_{t})\triangleq\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}\big{[}$ $1-\sum_{z_{t+1:{\tau}}}\sum_{x_{t:{\tau}}}b(x_{t})\overline{\mathbb{P}}(z_{t+1}\mid x_{t+1})\mathbb{P}(x_{t+1}\mid x_{t},a_{t})\cdot$ (11)
In the event that no subsequent observations are chosen for a given history, the value of $\bar{Q}^{\pi}(b_{t},a_{t})$ simplifies to the immediate reward plus an upper bound for any subsequent policy, given by $\mathcal{R}_{\max}\cdot(T-t-1)$.
Using UDB we define the action selection criteria according to
Moreover, the optimal value function can be bounded as follows,
Lemma 1

### v4

@@ 9 Mathematical Analysis
9.1 Theorem 1
Theorem 3.
@@ 9.1 Theorem 1
$\left|V^{\pi}(b_{t})\!-\!\bar{V}^{\pi}(b_{t})\right|\leq\!\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\!\sum_{z_{t+1:{\tau}}}\sum_{x_{t:{\tau}}}b(x_{t})\!\!\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]\triangleq\epsilon_{z}^{\pi}(b_{t}).$ (35)
Proof.
For notational convenience, we derive the bounds for the value function by denoting the prior belief as $b_{0}$,
$V_{0}^{\pi}(b_{0})=\mathbb{E}_{z_{1:T}}\left[\sum_{t=0}^{T}r(b_{t},a_{t})\right]$ (36)
applying the belief update equation,
$V_{0}^{\pi}(b_{0})$ $=\sum_{z_{1:T}}\prod_{\tau=1}^{T}\mathbb{P}\left(z_{\tau}\mid H_{\tau}^{-}\right)\sum_{t=0}^{T}\left[\sum_{x_{t}}\frac{\mathbb{P}(z_{t}\mid x_{t})\sum_{x_{t-1}}\mathbb{P}(x_{t}\mid x_{t-1},\pi_{t-1})b_{t-1}}{\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)}r(x_{t},a_{t})\right]$ (37)
$=\sum_{z_{1:T}}\prod_{\tau=1}^{T}\mathbb{P}\left(z_{\tau}\mid H_{\tau}^{-}\right)\sum_{t=0}^{T}\left[\sum_{x_{0:t}}\frac{\ \prod_{k=1}^{t}\ \mathbb{P}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})}{\prod_{\tau=1}^{t}\mathbb{P}\left(z_{\tau}\mid H_{\tau}^{-}\right)}r(x_{t},a_{t})\right]$ (38)
$=\sum_{t=0}^{T}\sum_{z_{1:T}}\sum_{x_{0:T}}\prod_{k=1}^{t}\ \mathbb{P}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})r(x_{t},a_{t})$ (39)
which applies similarly to the simplified value function,
$\bar{V}_{0}^{\pi}(b_{0})=\sum_{t=0}^{T}\sum_{z_{1:T}}\sum_{x_{0:T}}\prod_{k=1}^{t}\ \bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})r(x_{t},a_{t}).$ (40)
We begin the derivation by focusing on a single time step, $t$, and later generalize to the complete value function.
$|\mathbb{E}_{z_{1:t}}[r(b_{t})]-\bar{\mathbb{E}}_{z_{1:t}}[r(\bar{b}_{t})]|$ (41)
$=$ $|\sum_{z_{1:t}}\sum_{x_{0:t}}[\prod_{k=1}^{t}\ \mathbb{P}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})r(x_{t})$ (42)
$-\prod_{k^{\prime}=1}^{t}\ \bar{\mathbb{P}}(z_{k^{\prime}}\mid x_{k^{\prime}})\mathbb{P}(x_{k^{\prime}}\mid x_{k^{\prime}-1},\pi_{k^{\prime}-1})b(x_{0})r(x_{t})]|$
$\leq$ $\sum_{z_{1:t}}\sum_{x_{0:t}}\left|r(x_{t})\left[\prod_{k=1}^{t}\mathbb{P}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})-\prod_{k^{\prime}=1}^{t}b(x_{0})\ \bar{\mathbb{P}}(z_{k^{\prime}}\mid x_{k^{\prime}})\mathbb{P}(x_{k^{\prime}}\mid x_{k^{\prime}-1},\pi_{k^{\prime}-1})\right]\right|$ (43)
$=$ $\sum_{z_{1:t}}\sum_{x_{0:t}}|r(x_{t})|\left[\prod_{k=1}^{t}\mathbb{P}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})-\prod_{k^{\prime}=1}^{t}b(x_{0})\ \bar{\mathbb{P}}(z_{k^{\prime}}\mid x_{k^{\prime}})\ \mathbb{P}(x_{k^{\prime}}\mid x_{k^{\prime}-1},\pi_{k^{\prime}-1})\right]$ (44)
where the second transition is due to triangle inequality, the third transition is equality by the construction, i.e. using the simplified observation models imply that the difference is nonnegative. We add and subtract, followed by rearranging terms,
$=$ $\sum_{z_{1:t}}\sum_{x_{0:t}}|r(x_{t})|$ (45)
$[\prod_{k=1}^{t}\mathbb{P}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})-\prod_{k=1}^{t-1}b(x_{0})\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})$
$+\prod_{k=1}^{t-1}b(x_{0})\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})-\prod_{k^{\prime}=1}^{t}b(x_{0})\bar{\mathbb{P}}(z_{k^{\prime}},x_{k^{\prime}}\mid x_{k^{\prime}-1},\pi_{k^{\prime}-1})]$
[…]
$\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})\left[\prod_{k=1}^{t-1}\mathbb{P}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})-\prod_{k=1}^{t-1}b(x_{0})\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\right]$
$+\prod_{k=1}^{t-1}b(x_{0})\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})[\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})-\ \bar{\mathbb{P}}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})]\Bigr{\}}$
applying Holder’s inequality,
$\leq$ $\mathcal{R}_{\max}\sum_{z_{1:t}}\sum_{x_{0:t}}\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})\left[b(x_{0})\prod_{k=1}^{t-1}\mathbb{P}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})-b(x_{0})\prod_{k=1}^{t-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\right]$ (47)
$+\mathcal{R}_{\max}\sum_{z_{1:t}}\sum_{x_{0:t}}\prod_{k=1}^{t-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})[\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})-\ \bar{\mathbb{P}}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})]$
$=$ $\mathcal{R}_{\max}\sum_{z_{1:t}}\sum_{x_{0:t}}\mathbb{P}(z_{t},x_{t}\mid x_{t-1},\pi_{t-1})\cdot$ (48)
$\left[b(x_{0})\prod_{k=1}^{t-1}\mathbb{P}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})-b(x_{0})\prod_{k=1}^{t-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\right]+\mathcal{R}_{\max}\delta_{t}$
$=$ $\mathcal{R}_{\max}\sum_{z_{1:t-1}}\sum_{x_{0:t-1}}\left[b(x_{0})\prod_{k=1}^{t-1}\mathbb{P}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})-b(x_{0})\prod_{k=1}^{t-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})\right]$ (49)
$+\mathcal{R}_{\max}\delta_{t},$
following similar steps recursively,
$=\ldots=\mathcal{R}_{\max}\sum_{\tau=1}^{t}\delta_{\tau}.$ (50)
Finally, applying similar steps for every time step $t\in[1,T]$ results in,
$\left|V^{\pi}(b_{t})-\bar{V}^{\pi}(b_{t})\right|\leq\mathcal{R}_{\max}\sum_{t=1}^{T}\sum_{\tau=1}^{t}\delta_{\tau}$ (51)
where,
$\delta_{\tau}$ $=\sum_{z_{1:\tau}}\sum_{x_{0:\tau}}\prod_{k=1}^{\tau-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})[\mathbb{P}(z_{\tau},x_{\tau}\mid x_{\tau-1},\pi_{\tau-1})-\ \bar{\mathbb{P}}(z_{\tau},x_{\tau}\mid x_{\tau-1},\pi_{\tau-1})]$
$=\sum_{z_{1:\tau-1}}\sum_{x_{0:\tau-1}}\prod_{k=1}^{\tau-1}\bar{\mathbb{P}}(z_{k},x_{k}\mid x_{k-1},\pi_{k-1})b(x_{0})[1-\ \sum_{z_{\tau}}\sum_{x_{\tau}}\bar{\mathbb{P}}(z_{\tau},x_{\tau}\mid x_{\tau-1},\pi_{\tau-1})]$ (52)
plugging the term in (52) to (51) and expanding the terms results in the desired bound,
$\left|V^{\pi}(b_{t})-\bar{V}^{\pi}(b_{t})\right|\leq\!\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\!\sum_{z_{t+1:{\tau}}}\sum_{x_{t:{\tau}}}b(x_{t})\!\!\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]$ (53)
which concludes our derivation. ∎
9.2 Lemma 1
Lemma 4.